Carnot refrigeration cycle absorbs heat at 280k and rejects heat at 310k.
a: Calculate the coefficient of performance of this refrigeration cycle.
b: If the cycle is absorbing 1120KJ/min at 280k, how many KJ work is required per second .
C: If the carnot heat pump operates between the same temperature as the above refrigerator, what is the coefficient of performance ?
d: How many KJ/min will the heat pump deliver at 310k if it absorb 1120 KJ/min at 280k.
1
Expert's answer
2019-04-08T09:15:26-0400

a: Calculate the coefficient of performance of this refrigeration cycle.

Solution:

For Carnot refrigerator cycle:


"COP_R = \frac {T_L}{{T_H - T_L}} (1)"

Using (1) we got: COPR=9.34

Answer:

COPR=9.34


b: If the cycle is absorbing 1120KJ/min at 280k, how many KJ work is required per second

Solution:

The COP of a refrigerator is given by the following equation:


"COP_R=\frac {Q_L}{W_i} (2)"

Using (2) we got:




"W_i=\frac {Q_L}{COP_R} (3)"

"W_i=\frac {1120}{60\times 9.34}=2KJ"

Answer:

2KJ


C: If the carnot heat pump operates between the same temperature as the above refrigerator, what is the coefficient of performance ?


"COP_P=\frac {T_H}{T_H - T_L} (4)"

Using (4) we got: COPP=10.33


Answer:

10.33


d: How many KJ/min will the heat pump deliver at 310k if it absorb 1120 KJ/min at 280k.



"COP_P=\frac {Q_H}{W_i} (5)"

Using (5) we got:


"Q_H = COP_P \times W_i (6)"

QH= 11570 KJ/min

Answer:

11570 KJ/min